wiki:UC20180806

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 8/06/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Mary Beth Kaiser (JHU), John Bally (CU), Kal Kadlec (UW), Ben Williams (UW), Moire Prescott (NMSU), Bill Ketzeback (APO), Misty Bentz (GSU), Eric Bellm (UW), John Wisniewski (OU), Chip Kobulnicky (UWy), John Wilson (UVa), Sarah Tuttle (UW)


User feedback and comments from institutional representatives:

  • Colorado - nothing to report
  • Georgia State - nothing to report
  • JHU - nothing to report
  • NAPG - no report
  • NMSU - nothing to report, just a question on whether there is any news regarding the ARCTIC dark current issue that was raised in June? We will come back to this later in the meeting.
  • Oklahoma - nothing to report
  • Seattle - no report
  • UVa - no report
  • Washington - nothing to report
  • Wyoming - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 5/31/18 through 7/31/18

0) Overview

Monsoons started early this year so June observing was impacted a little more than nominal. Shutdown started on time and the small mirrors were delivered to the vendors on time. They are now at the vendors in various states of work.

1) Telescope

Telescope was operational until shutdown. Shutdown started July 5th and continues. M2 and M3 were removed and M1 was washed. M2 and M3 limit switches wholesale replacement is in work.

2) Instruments

DIS: adjusted collimation and camera servicing is in progress.

TripleSpec has been operational.

Agile has been operational this period.

Echelle is warm and on vacuum pump.

NIC-FPS is operational.

ARCTIC is operational. Some minor filter wheel work has been accomplished to address repeatability but it more controller work may be necessary.

3) ARCSAT

New guider scope has been installed and waiting for more engineering time to adjust and commission.


Additional telescope and instrument discussion:

Much progress has been made on regular servicing of the instruments and telescope itself. Kal was on site for several weeks and a fair amount of work was accomplished on recollimating DIS and understanding the problems with the scatter. DIS-Red had a bad batch of o-rings so is still in need of some additional work. Other instruments are being serviced as we speak. Some work was done on the ARCTIC filter wheel to make it more stable over the long term. We found some issues and we think we know how it should be fixed (needs different kind of programming using interrupts), but how to get there is still being discussed. Prior to shutdown everything was pretty normal.


ARCSAT Update:

The telescope is back together. We added a new guider and finder scope to the telescope that we're still in the process of commissioning. Based on weather and availability of personnel this may not get completed until later in the month.


Status of Secondary and Tertiary Mirrors:

At the start of summer shutdown M3 was sent out for polishing and recoating. It was decided that this needed to be done because the cleaning process we have been using seems to have caused acid etching on these particular mirrors. We saw it on the ARCSAT primary and M3, and a small amount on M2. To avoid making it worse we decided to polish it out of M3 and coat it with a protective UV enhanced coating. M3 was delivered to the polisher (in Tucson) on July 9. M2 is at the coaters (in California), along with the SDSS secondary mirror. The coaters need to build special tooling to hold the mirrors in place in the chamber, then they will do the coatings. M3 is at polisher and they’ve had a problem with their measuring machine and need to get that fixed before we can go any farther. That is an additional slip of the schedule but we do not know the full impact. After polishing, M3 will go to the coater and then it gets coated. Per a telecon earlier today with the Directors and observatory staff, we are progressing down this path. We looked at several alternatives but none proved to be worth pursuing, although we are still considering other options. Due to this schedule slip there will be a delay in the 3.5m return to science; it will probably be in mid-September but we don't have enough details yet to provide a firm date. The original estimated shutdown length was based on best-case parallel schedules of what the vendors gave us but we’re not in that best case, and we will fall out of their worst-case estimates at this point. Meanwhile we are doing work on the instruments and other telescope-related tasks.


Q3 3.5-m scheduling:

Users Committee reps were advised that if they have users who are scheduled on the 3.5m between 7-14 September there is very little chance that they will be able to conduct there observing. For users who are scheduled after 9/14 please warn them that even this is not guaranteed at this point. Russet intentionally designed the September schedule to have resilient programs scheduled right after shut down. Individual programs who lose time can talk to Ben, Nancy and Russet about trades with Q4 time. At this point the plan is to jump back into the schedule as soon as science resumes rather than trying to shuffle everyone around. Nancy agreed to provide email updates regarding the scheduled return to science to the Users Committee and Institutional Schedulers as they become available.


Q4 3.5-m scheduling:

Ben will send out the preliminary Q4 allocations to the Institutional Schedulers today or tomorrow.


Q4 ARCSAT scheduling:

The call for Q4 proposals for ARCSAT will come out around mid-September.


ARCTIC dark current issue:

As far as Bill knows no one is looking at ARCTIC dark current issue over the shutdown, unless someone is independently looking at old data. As a reminder, the issue raised by an NMSU user was that they noticed a lot of variation in dark current from exposure to exposure. Sarah asked and Bill confirmed that they detector temperature is regulated with a heater (yes). John Bally reported that he recently took a large series of darks, including some with very long (~ 1 hr) exposures, during some cloudy weather; the data are available if anyone wants to look at them. John Wisniewski requested clarification on whether this is an issue that affects short exposures or long, as this will impact different types of science depending on the answer. Nancy will check with Russet on the status of this issue.


NA2 Baffle Project

John Wilson provided a status report of this project. No progress was made over the summer, but as of late spring they had a full SolidWorks model of the telescope that best matches the Zemax prescription of telescope. Both of these pieces will be important for the last step, which will be engaging a firm that does scattered light analysis. John is now looking into how Zemax can be used to trace rays backwards from detector (really the focal plane of the telescope) in order to understand what the telescope sees (e.g. portions of the baffle). He hopes to make progress on this by mid-fall so that we have something to act upon.


KOSMOS update

Sarah Tuttle has received KOSMOS from KPNO! She is in the process of getting everything unpacked and turned on. They will do some basic quality checks before moving forward with installing the slit viewer and calibration lamps. She will update everyone as its capabilities become available. It is not likely to be ready for commissioning on the 3.5m until Q1 2019.


TEWG Update

The Training and Education Working Group (TEWG) has been active. They are nearing the end of their data-gathering phase and starting to think about how to synthesize the information learned. They are finding that different types of student groups have different requirements in terms of preparations and activities on site. A fair amount of discussion has focused on the scheduling of student trips to the site. Given the scheduling challenges that we are likely to face for Q4 given the extension of the summer shutdown in Q3, Nancy, Ben, and Russet ask that users who are planning to propose for a class trip should give them a heads up well before the proposal submission deadline.


Response to Call for White Papers for 2020 Decadal Survey

Nancy introduced this topic for discussion. The relevance/niche for 3-4m class telescopes should be addressed in the upcoming astrophysics decadal survey as they play an important role in certain types of science investigations, student training, instrument development and testing, etc. She reached out to directors of several other 3-6m telescopes and got a few positive responses. Questions for discussion were: Should we do this? Should we try to collaborate with others? Do we want to cover the broader landscape of 1-6m telescopes or keep it more focused? The general consensus was that yes we should do it (and ideally LEAD this effort), and yes we should try to collaborate with others because it sends a more powerful message that way. Suggested way forward: collect the science cases for things we already know we're doing with the 3.5m and get a clearer focus for the white paper, then seek specific input from broader user community if needed.


Scheduling Our Meetings

The times for the fall semester Users Committee meetings will be at 10:30 am Mountain Time. Whereas in the past we have tended to meet on the first Monday of each month, this fall the dates will move around a bit to avoid several holidays and the ARC Board of Governors meeting. Thus, the dates for our meetings will be as follows:

  • Wednesday September 5
  • Monday October 1
  • Monday October 29
  • Monday December 3

All Other Business

John Bally asked if anyone has a distortion map for ARCTIC. This is needed if one wants to make mosaics from a bunch of images. Nancy suggested that he should check with Joanne Hughes Clark - Nancy will connect the two of them by email.


ACTION ITEMS:

Action items from previous meetings:

None.

New action items from this meeting:

  • Nancy will provide email updates to the Users Committee and Institutional Schedulers when there are updates on the scheduled return to science.
  • Users Committee reps are asked to notify their users who are scheduled to observe in September of the summer shutdown schedule slip.
  • Users Committee reps are asked to communicate to their faculty that if anyone is planning to propose a class training trip to the site for Q4 2018 they should notify Nancy, Ben, and Russet ASAP (before the proposal deadline!).
  • Nancy will check with Russet on the status of the ARCTIC dark current issue.
  • Nancy will connect John Bally and Joanne Hughes Clark re: ARCTIC distortion maps. STATUS: CLOSED. Introduction email sent 8/6/18. Email discussion among several interested parties is ongoing.

Next meeting:

The next Users Committee telecon will be on Wednesday September 5 at 10:30 am MDT.

Note: See TracWiki for help on using the wiki.